Find two numbers such that one number is 5 greater than the other number. If the sum of their squares is 5 times the square of half the smaller number, what are the numbers?
Answer by lwsshak3(11628)

Find two numbers such that one number is 5 greater than the other number. If the sum of their squares is 5 times the square of half the smaller number, what are the numbers?
***
let x=smaller number
x+5=larger number
..
x^2+(x+5)^2=5x^2
x^2+x^2+10x+25=5x^2
3x^2-10x-25=0
(3x+5)(x-5)=0
x=-5/3
or
x=5
x+5=10
smaller number=5
larger number=10